y Q = 300√ L − 4L, where Q denotes output and L denotes the size of workfor e, al ulate the value of marginal produ t of labour if L = 9


1
Expert's answer
2021-07-12T13:11:01-0400

Solution.


"Q=300\sqrt{l}-4l."


Marginal product

"MP=\frac{dQ}{dl}."

"MP=300\frac{1}{2\sqrt{l}}-4=\frac{150}{\sqrt{l}}-4."

If "l=9," then

"MP=\frac{150}{\sqrt{9}}-4=46."

Answer. 46.